EXTENSION OF THE STDS/HTDS SOFTWARE TO ROVIBRONIC SPECTROSCOPY AND TO LOWER SYMMETRY PROBLEMS

\underline{V. BOUDON}, M. REY, M. ROTGER, M. LOETE, Laboratoire de Physique de l’Université de Bourgogne, CNRS UMR 5027, 9 Av. A. Savary, BP 4780, F-21078 DIJON Cedex, FRANCE. E-mail : boudon@jupiter.u-bourgogne.fr.

We present new softwares using the group-theoretical and tensorial methods developed in the Dijon group. They are based on the Spherical Top Data System (STDS)^{a} and the Highly-spherical Top Data System (HTDS)^{b} previously written for spherical-top molecules in a singlet electronic state. The two types of problems that can be handled with these new programs are :

• The case of spherical-top molecules in a degenerate electronic state. This concerns the rovibronic spectroscopy instead of the rovibrational spectroscopy, for species like ReF_{6} or V(CO)_{6}. A new formalism including electronic operators and rovibronic couplings has been developed^{c}^{,}^{d} and implemented in a new version of HTDS.

• The case of “quasi spherical-top” molecules of C_{4v} and C_{2v} symmetry. This concerns symmetric-top molecules deriving from spherical-tops by substitution of one ligand like SF_{5}Cl (C_{4}v symmetry) or asymmetric-tops with ligands having close masses like SO_{2}F_{2} (C_{2}v symmetry). New formalism using the O(3) ⊃ O_{h} ⊃ C_{4v}^{e}^{,}^{f} and O(3) ⊃ T_{d} ⊃ C_{2v} group chains have been developed and implemented into two new program suites named C_{4v}TDS and C_{2v}TDS, respectively.

These software packages will be available soon for free download.

aCh. Wenger and J.-P. Champion, J. Quant. Spectrosc. Radiat. Transfer, 59, 471–480 (1998).

bCh. Wenger, V. Boudon J.-P. Champion and G. Pierre, J. Quant. Spectrosc. Radiat. Transfer, 66, 1–16 (2000).

cM. Rey, V. Boudon, M. Loëte and F. Michelot, J. Mol. Spectrosc., 204, 106–119 (2000).

dM. Rey, V. Boudon and M. Loëte, J. Mol. Struct., in press (2001).

eM. Rotger, V. Boudon and M. Loëte, J. Mol. Spectrosc., 200, 123–130 (2000).

fM. Rotger, V. Boudon and M. Loëte, J. Mol. Spectrosc., 200, 131–137 (2000).